Show that on the curve y = 3x2 − 7x + 6 the chord joining the points whose abscissa

are x = 1 and x = 2, is parallel to the tangent at the whose abscissa is x= 3/2



1
Expert's answer
2021-06-13T17:29:44-0400

Solution

Ordinates for two points of chord abscissa x1=1 and x2=2 are y1=y(x1)=2 and y2= y(x2)=4.

The slope of the chord is k=(y2-y1)/(x2-x1)=2.

The tangent at the any point of curve y(x) is the line with the slope equal to its derivative y’(x)=6x-7. At the abscissa x0=3/2 the tangent is y’(3/2)=2.

So slopes of the chord and the tangent are equal. Therefore this lines are parallel.